# A Quasar-Anchored Protocluster at in the ASPIRE Survey: II. An Environmental Analysis of Galaxy Properties in an Overdense Structure

### v1

@@ 1 Introduction
Quasars hosting supermassive black holes (SMBH) with masses exceeding M⊙ are now routinely observed in the high-redshift () Universe (Fan et al., 2023), but challenges still exist in identifying the mechanisms responsible for their formation and growth. Because they reside in massive dark matter halos (e.g., Costa et al., 2014; Zhang et al., 2023), they are theoretically expected to trace large-scale galaxy overdensities, but observations have reached little consensus on whether this is routinely the case (e.g., Kim et al., 2009; Bañados et al., 2013; Mazzucchelli et al., 2017; Ota et al., 2018; Champagne et al., 2023; Lambert et al., 2024; Rojas-Ruiz et al., 2024). While previous observations were typically limited to photometric identification of companions and/or observations in small fields of view, significant variance among quasar environments is seen even with the best observing mode available — spectroscopy in wide fields, now available with JWST mosaics.
@@
Two reionization-era quasar surveys in particular, EIGER (Matthee et al., 2023a; Kashino et al., 2022; Eilers et al., 2024) and ASPIRE (Wang et al., 2023; Yang et al., 2023) have undertaken grism spectroscopy with JWST/NIRCam to identify H+[Oiii]-emitters at in the fields of a total of 31 quasars at . Several of these fields contain notable () overdensities of companion galaxies (Wang et al. in prep, Champagne et al. 2024a, Wang et al., 2023; Matthee et al., 2023a), but we are still lacking detailed case studies investigating these putative protoclusters. A number of protocluster candidates at identified as serendipitous overdensities in survey fields have been spectroscopically confirmed with JWST and investigated in detail (e.g., Morishita et al., 2023; Helton et al., 2023; Sun et al., 2023), but targeted studies of protoclusters anchored by quasars are absent from these samples.
@@
To this end, this paper focuses on the quasar field J03053150, which has been exhaustively studied from optical to submm wavelengths (Venemans et al., 2013; Farina et al., 2017; Champagne et al., 2018; Mazzucchelli et al., 2017; Venemans et al., 2019; Meyer et al., 2022; Wang et al., 2023, Champagne et al. 2024a). J03053150 was originally identified in the VIKING Survey (Venemans et al., 2013, 2016) and lies at , containing a M⊙ SMBH. Previously known to host LBG and [Cii]-emitter overdensities (Ota et al., 2018; Venemans et al., 2019; Champagne et al., 2023), then Wang et al. (2023) used Cycle 1 ASPIRE data to identify 41 galaxies at in a single NIRCam/WFSS pointing of the same field via the detection of [Oiii]5007 — given that 21 of these were within from the quasar, the evidence suggested the existence of a protocluster.
@@
Because protoclusters at are expected to span several tens of comoving Mpc (e.g., Chiang et al., 2017), Paper I of this series (Champagne et al. 2024a) presented Cycle 2 data consisting of a 35 arcmin2 NIRCam/WFSS mosaic around J03053150 to search for additional members of the putative protocluster. We found a total of 124 [Oiii] emitters in the expanded mosaic, with a remarkable 53 members of the structure spanning the full NIRCam footprint ( Mpc) and centered on the quasar redshift. The richness and spatial clustering of the overdensity suggests that it is a cluster progenitor similar to those seen in simulations at least out to (e.g., Bassini et al., 2020; Remus et al., 2023), but observational definitions of “protoclusters” at high redshift are discrepant and mostly rely on identifications of overdensities of star-forming galaxies within some redshift and projected distance cut (e.g., Herard-Demanche et al., 2023; Helton et al., 2023; Morishita et al., 2023). Much is lacking in terms of our knowledge of stellar mass assembly and galaxy/black hole evolution within the earliest protoclusters. With spectroscopic data in hand and a suite of imaging from prior studies (Ota et al., 2018; Champagne et al., 2023) and NIRCam (Wang et al., 2023), we can investigate in detail the rest-optical properties of these galaxies and compare them to the general field population towards the end of the epoch of reionization (EoR).
@@ 2.1 JWST Data
J03053150 was observed as part of the Cycle 1 JWST ASPIRE (A SPectroscopic survey of biased halos In the Reionization Era) program (GO #2078, PI: F. Wang) which targets 25 quasars with F356W grism spectroscopy and F115W/F200W/F356W broadband imaging with NIRCam. More details about the ASPIRE survey and its Cycle 2 follow-up observations can be found in Wang et al. (2023) and Champagne et al. 2024a (hereafter Paper I). Followup mosaic observations with the same setup were performed in Cycle 2 (GO #3325, PI: F. Wang), constituting 5 additional pointings covering a total area of 35.09 arcmin2. Data reduction, background subtraction, and astrometric alignment of the images are all described in Paper I. That paper also describes the spectral extraction procedure that produces our final sample of 124 line emitters. To better constrain the SEDs of our target galaxies, we also include archival imaging from HST and Subaru.
@@ 4 SED fitting
Between the two codes, we find that Bagpipes returns, on average, stellar masses 0.25 dex higher, SFRs 0.44 dex higher, and ages of 0.07 dex younger than that with Prospector; we note that there is a strong degeneracy between the age of the SED and the reported stellar mass as has been noted by other authors (e.g., Whitler et al., 2022; Endsley et al., 2021). We also find that Bagpipes and Prospector tend to disagree in the shape of the rest-optical continuum due to our poor wavelength coverage beyond rest-frame 0.5 m; particularly, Prospector fits a number of large Balmer jumps leading to very young ages for the galaxies, but we cannot necessarily rule these solutions out. In the following analysis, we use Prospector as our fiducial SED model as it typically returns lower values for our entire galaxy sample, and note the differences in various trends when using Bagpipes where relevant. We use the median value of the posterior distributions of each galaxy property as the reported value, and the 16th and 84th percentiles as our 1 errors.
@@ 5.2 Occurrence Rate of AGN
To check whether any of our sources were AGN, we performed three tests: 1) searching for broadened H emission (e.g. Matthee et al., 2023b; Greene et al., 2023) 2) using the mass-excitation diagram, i.e. log([Oiii]/H) versus stellar mass (Juneau et al., 2011, 2014), and 3) using “little red dot” photometric criteria (e.g., Greene et al., 2023; Kokorev et al., 2024; Akins et al., 2024). However, we note that our current data is limited in discriminatory power in determining the presence of AGN. We detect H in roughly half of the sample, and we use 3 upper limits for 65/124 galaxies. We also do not have the wavelength coverage to access the high-ionization lines that would support the presence of AGN such as [N iv]1487, C iv1549, or He ii1640 (e.g., Scholtz et al., 2023). Finally, the “little red dot” phenomenon (e.g., Matthee et al., 2023b; Greene et al., 2023; Harikane et al., 2023; Kokorev et al., 2024; Akins et al., 2024) characterized by a steeply rising red optical continuum suspected to be due to AGN, is not easily isolated without a photometric band clean of line contribution (e.g, F277W or F444W).
@@
Next, we can compare the flux ratio of [Oiii]/H to stellar mass (i.e., the mass-excitation diagram or MEx; Juneau et al., 2011) to infer whether the line emission is powered primarily by star formation or an AGN. We present this in Figure 7 and compare to the demarcation between SF and AGN from Juneau et al. (2014). We find that 35/124 galaxies in the full sample are consistent with lying above the star formation boundary derived by Juneau et al. (2014), though we note we are comparing to an AGN/SF boundary calibrated at which might not hold for the typically lower-metallicity galaxies in the early Universe (e.g., Shapley et al., 2023). Indeed, both observational (e.g., Scholtz et al., 2023; Matthee et al., 2023a) and theoretical (e.g., Hirschmann et al., 2023) results indicate that the ratio of [Oiii] to H at fixed stellar mass increases with redshift due to generally higher ionization parameters fueling higher specific SFR and, secondarily, lower metallicity within the ISM of high redshift galaxies, especially pronounced at lower stellar masses. In Figure 7 we also show the expected mass-excitation relationship based on the strong-line calibrations derived by Bian et al. (2018) for nearby high- analogues and the mass-metallicity relationship (MZR) at from the FIRE simulations in Ma et al. (2016)333The MZR relationship is scaled to the EIGER stacked observations as in Matthee et al. (2023a).: though there is substantial scatter, our data is consistent with the shape of the theoretically-derived MZR and the stacked observational results from EIGER at (Matthee et al., 2023a). While 14/35 [Oiii]-emitters identified in the MEx diagram above the SF boundary lie within the protocluster, we are not confident that any of these galaxies are AGN based on this selection alone.
Finally, we identify “little red dots” (LRDs; e.g., Matthee et al., 2023b; Greene et al., 2023) which are characterized by blue UV slopes and red optical slopes (only some of which show broadened Balmer lines; see Harikane et al., 2023), suggesting contribution of an AGN to the galaxy’s total light output. To perform this search, we modify the color selection and compactness criteria from Greene et al. (2023) to match our filter coverage. Specifically, we impose the following color criteria444While having data in F444W would be ideal to probe the rest-optical continuum, this criteria set still searches for a red continuum slope and filters out the colors of brown dwarfs in the galaxy (which is not a concern since we only perform this search among confirmed line emitters).:

### v2

@@ 1 Introduction
Quasars hosting supermassive black holes (SMBH) with masses exceeding M⊙ are now routinely observed in the high-redshift () Universe (Fan et al., 2023), but challenges still exist in identifying the mechanisms responsible for their formation and growth. Because they reside in massive dark matter halos (e.g., Costa et al., 2014; Zhang et al., 2024), they are theoretically expected to trace large-scale galaxy overdensities, but observations have reached little consensus on whether this is routinely the case (e.g., Kim et al., 2009; Bañados et al., 2013; Mazzucchelli et al., 2017; Ota et al., 2018; Champagne et al., 2023; Lambert et al., 2024; Rojas-Ruiz2024a). While previous observations were typically limited to photometric identification of companions and/or observations in small fields of view, significant variance among quasar environments is seen even with the best observing mode available — spectroscopy in wide fields, now available with JWST mosaics.
@@
Two reionization-era quasar surveys in particular, EIGER (Matthee et al., 2023; Kashino et al., 2023; Eilers et al., 2024) and ASPIRE (Wang et al., 2023; Yang et al., 2023) have undertaken grism spectroscopy with JWST/NIRCam to identify H+[Oiii]-emitters at in the fields of a total of 31 quasars at . Several of these fields contain notable () overdensities of companion galaxies (Wang et al. in prep, Wang et al., 2023; Matthee et al., 2023; Champagne et al., 2024), but we are still lacking detailed case studies investigating these putative protoclusters. A number of protocluster candidates at identified as serendipitous overdensities in survey fields have been spectroscopically confirmed with JWST and investigated in detail (e.g., Morishita et al., 2023; Sun et al., 2023; Helton et al., 2024), but targeted studies of protoclusters anchored by quasars are absent from these samples.
@@
To this end, this paper focuses on the quasar field J03053150, which has been exhaustively studied from optical to submm wavelengths (Venemans et al., 2013; Farina et al., 2017; Champagne et al., 2018; Mazzucchelli et al., 2017; Venemans et al., 2019; Meyer et al., 2022; Wang et al., 2023; Champagne et al., 2024). J03053150 was originally identified in the VIKING Survey (Venemans et al., 2013, 2016) and lies at , containing a M⊙ SMBH. Previously known to host LBG and [Cii]-emitter overdensities (Ota et al., 2018; Venemans et al., 2019; Champagne et al., 2023), then Wang et al. (2023) used Cycle 1 ASPIRE data to identify 41 galaxies at in a single NIRCam/WFSS pointing of the same field via the detection of [Oiii]5007 — given that 21 of these were within from the quasar, the evidence suggested the existence of a protocluster.
@@
Because protoclusters at are expected to span several tens of comoving Mpc (e.g., Chiang et al., 2017), Paper I of this series (Champagne et al., 2024) presented Cycle 2 data consisting of a 35 arcmin2 NIRCam/WFSS mosaic around J03053150 to search for additional members of the putative protocluster. We found a total of 124 [Oiii] emitters in the expanded mosaic, with a remarkable 53 members of the structure spanning the full NIRCam footprint ( Mpc) and centered on the quasar redshift. The richness and spatial clustering of the overdensity suggests that it is a cluster progenitor similar to those seen in simulations at least out to (e.g., Bassini et al., 2020; Remus et al., 2023), but observational definitions of “protoclusters” at high redshift are discrepant and mostly rely on identifications of overdensities of star-forming galaxies within some redshift and projected distance cut (e.g., Morishita et al., 2023; Helton et al., 2024; Herard-Demanche et al., 2025). Much is lacking in terms of our knowledge of stellar mass assembly and galaxy/black hole evolution within the earliest protoclusters. With spectroscopic data in hand and a suite of imaging from prior studies (Ota et al., 2018; Champagne et al., 2023) and NIRCam (Wang et al., 2023), we can investigate in detail the rest-optical properties of these galaxies and compare them to the general field population towards the end of the epoch of reionization (EoR).
@@ 2.1 JWST Data
J03053150 was observed as part of the Cycle 1 JWST ASPIRE (A SPectroscopic survey of biased halos In the Reionization Era) program (GO #2078, PI: F. Wang) which targets 25 quasars with F356W grism spectroscopy and F115W/F200W/F356W broadband imaging with NIRCam. More details about the ASPIRE survey and its Cycle 2 follow-up observations can be found in Wang et al. (2023) and Champagne et al. (2024), hereafter Paper I. Followup mosaic observations with the same setup were performed in Cycle 2 (GO #3325, PI: F. Wang), constituting 5 additional pointings covering a total area of 35.09 arcmin2. Data reduction, background subtraction, and astrometric alignment of the images are all described in Paper I. That paper also describes the spectral extraction procedure that produces our final sample of 124 line emitters. To better constrain the SEDs of our target galaxies, we also include archival imaging from HST and Subaru.
@@ 4 SED fitting
Between the two codes, we find that Bagpipes returns, on average, stellar masses 0.25 dex higher, SFRs 0.44 dex higher, and ages of 0.07 dex younger than that with Prospector; we note that there is a strong degeneracy between the age of the SED and the reported stellar mass as has been noted by other authors (e.g., Endsley et al., 2021; Whitler et al., 2023). We also find that Bagpipes and Prospector tend to disagree in the shape of the rest-optical continuum due to our poor wavelength coverage beyond rest-frame 0.5 m; particularly, Prospector fits a number of large Balmer jumps leading to very young ages for the galaxies, but we cannot necessarily rule these solutions out. In the following analysis, we use Prospector as our fiducial SED model as it typically returns lower values for our entire galaxy sample, and note the differences in various trends when using Bagpipes where relevant. We use the median value of the posterior distributions of each galaxy property as the reported value, and the 16th and 84th percentiles as our 1 errors.
@@ 5.2 Occurrence Rate of AGN
To check whether any of our sources were AGN, we performed three tests: 1) searching for broadened H emission (e.g. Matthee et al., 2024; Greene et al., 2024) 2) using the mass-excitation diagram, i.e. log([Oiii]/H) versus stellar mass (Juneau et al., 2011, 2014), and 3) using “little red dot” photometric criteria (e.g., Greene et al., 2024; Kokorev et al., 2024; Akins et al., 2024). However, we note that our current data is limited in discriminatory power in determining the presence of AGN. We detect H in roughly half of the sample, and we use 3 upper limits for 65/124 galaxies. We also do not have the wavelength coverage to access the high-ionization lines that would support the presence of AGN such as [N iv]1487, C iv1549, or He ii1640 (e.g., Scholtz et al., 2023). Finally, the “little red dot” phenomenon (e.g., Matthee et al., 2024; Harikane et al., 2023; Greene et al., 2024; Kokorev et al., 2024; Akins et al., 2024) characterized by a steeply rising red optical continuum suspected to be due to AGN, is not easily isolated without a photometric band clean of line contribution (e.g, F277W or F444W).
@@
Next, we can compare the flux ratio of [Oiii]/H to stellar mass (i.e., the mass-excitation diagram or MEx; Juneau et al., 2011) to infer whether the line emission is powered primarily by star formation or an AGN. We present this in Figure 7 and compare to the demarcation between SF and AGN from Juneau et al. (2014). We find that 35/124 galaxies in the full sample are consistent with lying above the star formation boundary derived by Juneau et al. (2014), though we note we are comparing to an AGN/SF boundary calibrated at which might not hold for the typically lower-metallicity galaxies in the early Universe (e.g., Shapley et al., 2023). Indeed, both observational (e.g., Scholtz et al., 2023; Matthee et al., 2023) and theoretical (e.g., Hirschmann et al., 2023) results indicate that the ratio of [Oiii] to H at fixed stellar mass increases with redshift due to generally higher ionization parameters fueling higher specific SFR and, secondarily, lower metallicity within the ISM of high redshift galaxies, especially pronounced at lower stellar masses. In Figure 7 we also show the expected mass-excitation relationship based on the strong-line calibrations derived by Bian et al. (2018) for nearby high- analogues and the mass-metallicity relationship (MZR) at from the FIRE simulations in Ma et al. (2016)333The MZR relationship is scaled to the EIGER stacked observations as in Matthee et al. (2023).: though there is substantial scatter, our data is consistent with the shape of the theoretically-derived MZR and the stacked observational results from EIGER at (Matthee et al., 2023). While 14/35 [Oiii]-emitters identified in the MEx diagram above the SF boundary lie within the protocluster, we are not confident that any of these galaxies are AGN based on this selection alone.
Finally, we identify “little red dots” (LRDs; e.g., Matthee et al., 2024; Greene et al., 2024) which are characterized by blue UV slopes and red optical slopes (only some of which show broadened Balmer lines; see Harikane et al., 2023), suggesting contribution of an AGN to the galaxy’s total light output. To perform this search, we modify the color selection and compactness criteria from Greene et al. (2024) to match our filter coverage. Specifically, we impose the following color criteria444While having data in F444W would be ideal to probe the rest-optical continuum, this criteria set still searches for a red continuum slope and filters out the colors of brown dwarfs in the galaxy (which is not a concern since we only perform this search among confirmed line emitters).: